# arxiv-appendix

### v1

@@ 4.2 Real Data Experiments
In this section, we further evaluate Algorithm 2 and 3 on four public classification datasets, whose summary statistics are shown in Table 1. We adopt the framework of creating bandit instances from existing classification datasets following Riquelme et al. (2018) and Dimakopoulou et al. (2019). Specifically, we now have observations $\{X_{i},\dot{Y}_{i}\}_{i=1}^{n_{\mathrm{P}}}$, where the label $\dot{Y}_{i}\in[K]$. Let the underlying relationship between $\dot{Y}$ and $X$ be $\dot{f}_{k}(X):=\mathbb{P}\left[\dot{Y}=k|X\right]$ for $k\in[K]$. We regard the $K$ classes as bandit arms. We return a reward $r_{k,\ell}$ if $\dot{Y}=k$ and the policy pulls the arm $\ell$, i.e. $\pi(X)=\ell$, where $k,\ell\in[K]$. Here, $r_{k,\ell}$ can be either a constant or a random variable. The evaluation metric is then $\sum_{i=1}^{n_{\mathrm{P}}}r_{\dot{Y}_{i},\pi_{i}(X_{i})}$, where larger is better.
Such transformation, though wildly adopted, has never been explained statistically. We now show this is reasonable. In this framework, randomness comes from the noisy label $\dot{Y}$ instead of calculating $r_{\dot{Y}_{i},\pi_{i}(X_{i})}$ itself. The reward function is then
$f_{\ell}(x):=\mathbb{E}\left[r_{\dot{Y}_{i},\ell}\right]=\sum_{k=1}^{K}\dot{f}_{k}(x)\mathbb{E}_{r}\left[r_{k,\ell}\right],\;\;\ell\in[K],$
which is a weighted sum of the original class probability functions. In situations where it is appropriate to assume the class probability functions are smooth, $f_{\ell}$ is also smooth. The expectation of the evaluation metric has
$\mathbb{E}_{X,\dot{Y},r}\left[\sum_{i=1}^{n_{\mathrm{P}}}r_{\dot{Y}_{i},\pi_{i}(X_{i})}\right]=$ $\sum_{i=1}^{n_{\mathrm{P}}}\mathbb{E}_{X,r}\left[\sum_{k=1}^{K}\dot{f}_{k}(X_{i})r_{k,\pi_{i}(X_{i})}\right]$
$=$ $\sum_{i=1}^{n_{\mathrm{P}}}\mathbb{E}_{X}\left[f_{\pi_{i}(X_{i})}(X_{i})\right]$
which is compatible with (1). We collect the detailed information and pre-processing steps for each dataset in Appendix LABEL:sec:datasetdetails, as well as the choice of $r_{k,l}$.
Comparison methods considered are:
@@
  • LDPMAB: the proposed method for locally private contextual multi-armed bandits. We test with and without (marked as w and wo, respectively) auxiliary data.
@@
  • Linear: the generalized linear bandits with LDP proposed by Han et al. (2021). We choose (multinomial) logistic regression. We also test the method with auxiliary datasets, sequentially including auxiliary data directly for stochastic gradient descent with the required privacy level.
@@
  • NN: we generalize Linear by replacing the model with a single-layer neural network, with the other steps remaining unchanged.
The implementation details are deferred to Appendix LABEL:sec:implementationdetail. We first compute the reward across 30 random experiments, with randomness from partitioning (e.g. variability in the max-edge rule), the privacy mechanism, different initializations (e.g. initial parameter values in Linear and NN), random shuffling, and source policy realizations. To standardize the scale across datasets, we report the ratio of mean reward relative to that of non-privately fitting over target samples using ABSE in Perchet and Rigollet (2013). Thus, a reported value larger than $1$ means that the method is better than ABSE on only target data, and vice versa. The results are shown in Table 2.
@@
Several observations are in order. First, the proposed method with auxiliary data outperforms its competitors in terms of both best performance (number of significantly better rewards) and average performance (rank-sum). This shows that the proposed methods can effectively utilize auxiliary data. In contrast, Linear and NN occasionally have negative transfer, where auxiliary data actually worsen the performance. Moreover, the proposed method without auxiliary data shows comparable performances to the other methods that apply only target data. The Jester dataset shows a strong linearity nature, where Linear based methods significantly outperform the other non-linear approaches. Compared to ABSE, the competing methods are usually worse since privacy is applied.
@@
Table 2: The best results are marked as bold. To ensure significance, we adopt the Wilcoxon signed-rank test (Wilcoxon, 1992) with a significance level of 0.05 to check if the result is significantly better. The best results that hold significance towards the others are highlighted in grey. Due to the large sizes of the datasets, we report the performance at both $t=n_{\mathrm{P}}/4$ and $t=n_{\mathrm{P}}$ to highlight distinctions. We choose $\varepsilon\in\{1,2\}$ and $\varepsilon_{m}\in\{1,4\}$.
@@ 5 Conclusions and Discussions
In this work, we investigate the problem of contextual multi-armed bandits under local differential privacy. Under standard assumptions on nonparametric reward functions, we propose a near-optimal uniform-confidence-bound-type estimator, with its near-optimality supported by a minimax lower bound. To further improve the limited ceiling of LDP contextual MAB, we incorporate side information from auxiliary datasets that are also subject to LDP constraints. Assuming covariate shift, we introduce a jump-start scheme to leverage the auxiliary data, attaining the established minimax lower bound, up to logarithmic factors in interesting regimes. Extensive experiments on synthetic and real datasets validate our theoretical findings and demonstrate the superiority of our methodology.
We remark on the implications of our method in the context of multi-task learning. Consider a scenario where a set of $M$ players are deployed to engage in a bandit game, with the overall objective being to minimize the average regret across all players (Deshmukh et al., 2017; Wang et al., 2021). These players simultaneously interact with a shared set of arms. At each round, each player selects an arm and receives feedback. Each arm’s reward distribution is identical across all players. The estimator in (26) is permutation invariant with respect to the datasets. This means that treating any dataset as the target dataset does not affect the estimator’s effectiveness or the subsequent confidence bound (28). This observation suggests that the proposed methodology can be extended to multi-task learning, provided Algorithm 3 is adapted to accommodate parallel interactions. However, it is crucial to note that the assumption of identical reward distributions is fundamental, as it ensures the permutation invariance between the distributions.

### v2

@@ 4.2 Real Data Experiments
In this section, we further examine the performance of the proposed algorithms on three widely used classification datasets, whose summary statistics are given in Table 1. The detailed information for each dataset, including covariates, responses, pre-processing and selection of target and auxiliary data, are collected in Section LABEL:sec:datasetdetails of the supplement.
@@
In particular, we adopt the framework of creating bandit instances from (offline) classification datasets following Riquelme et al., (2018) and Dimakopoulou et al., (2019). Suppose we have a classification dataset $\{X_{i},\dot{Y}_{i}\}_{i=1}^{n_{\mathrm{P}}}$, where the class labels $\dot{Y}_{i}\in[K]$. We regard the $K$ classes as the bandit arms and define the reward of the $k$-th arm as $Y_{i}^{(k)}=\mathbf{1}(\dot{Y}_{i}=k)$. Let the underlying true relationship between $\dot{Y}$ and $X$ be $\dot{f}_{k}(X):=\mathbb{P}\big{[}\dot{Y}=k|X\big{]}$ for $k\in[K]$. This implies that the expected reward function of the $k$-th arm can be computed as
$f_{k}(x):=\mathbb{E}\left[Y_{i}^{(k)}|X_{i}=x\right]=\dot{f}_{k}(x),\text{ for all }k\in[K].$
Thus, if the class probability functions are smooth, the reward function $f_{k}$ is also smooth.
The evaluation metric is defined as the cumulative reward $\sum_{i=1}^{n_{\mathrm{P}}}{Y}_{i}^{\pi_{i}(X_{i})}$, with an expectation
$\mathbb{E}_{X,\dot{Y}}\left[\sum_{i=1}^{n_{\mathrm{P}}}{Y}_{i}^{\pi_{i}(X_{i})}\right]=\sum_{i=1}^{n_{\mathrm{P}}}\mathbb{E}_{X}\left[\sum_{k=1}^{K}\dot{f}_{k}(X_{i})\mathbf{1}(\pi_{i}(X_{i})=k)\right]=\sum_{i=1}^{n_{\mathrm{P}}}\mathbb{E}_{X}\left[f_{\pi_{i}(X_{i})}(X_{i})\right],$
which is compatible with the regret defined in (1). Note that since the true class probability functions $\{\dot{f}_{k}(\cdot)\}$ are unknown, we cannot directly compute the reward as $\sum_{i=1}^{n_{\mathrm{P}}}f_{\pi_{i}(X_{i})}(X_{i})$.
We consider three competing methods and a benchmark method:
@@
  • LDPMAB: our proposed method for LDP contextual nonparametric multi-armed bandits. We implement LDPMAB with and without (marked as w and wo, respectively) auxiliary data.
@@
  • Linear: the method proposed in Han et al., (2021) for LDP contextual generalized linear bandits (see Algorithm 2 therein), which does not consider transfer learning. We set the parametric model for the expected reward of each arm as a logistic function. We also test the method with auxiliary data, where we include auxiliary data in the stochastic gradient descent of the parameter estimation with the required privacy level.
@@
  • NN: we generalize Linear by replacing the expected reward model for each arm with a single-layer neural network, with the other steps staying unchanged.
  • ABSE: the method proposed in Perchet and Rigollet, (2013) for non-private contextual nonparametric multi-armed bandits, which does not consider transfer learning.
The implementation details of all methods can be found in LABEL:sec:implementationdetail of the supplement and we present the experiment result based on 100 repetitions. To proceed, we first explain how the experiment is implemented for each repetition (for simplicity of presentation, we assume $M=1$). In particular, given the original target data $\{X_{i}^{\mathrm{P}},\dot{Y}_{i}^{\mathrm{P}}\}_{i=1}^{n_{\mathrm{P}}}$ and auxiliary data $\{X_{i}^{\mathrm{Q}},\dot{Y}_{i}^{\mathrm{Q}}\}_{i=1}^{n_{\mathrm{Q}}}$ from the (offline) classification dataset, the following steps are executed sequentially:
  • We first conduct a random permutation of the index $\{1,2,\cdots,n_{\mathrm{P}}\}$ and $\{1,2,\cdots,n_{\mathrm{Q}}\}$. With an abuse of notation, we denote the permuted data via $\{X_{i}^{\mathrm{P}},\dot{Y}_{i}^{\mathrm{P}}\}_{i=1}^{n_{\mathrm{P}}}$ and $\{X_{i},\dot{Y}_{i}^{\mathrm{Q}}\}_{i=1}^{n_{\mathrm{Q}}}$ as well.
  • We now generate the bandit auxiliary data. For each $i\in[n_{\mathrm{Q}}]$, given $X_{i}^{\mathrm{Q}}$, we implement the behavior policy $\pi^{\mathrm{Q}}$, pull arm $\pi^{\mathrm{Q}}(X_{i}^{\mathrm{Q}})$ and observe the reward $Y_{i}^{\mathrm{Q},(\pi^{\mathrm{Q}}(X_{i}^{\mathrm{Q}}))}:=\mathbf{1}(\dot{Y}_{i}^{\mathrm{Q}}=\pi^{\mathrm{Q}}(X_{i}^{\mathrm{Q}}))$. We thus attain the bandit auxiliary data $\mathcal{D}^{\mathrm{Q}}=\{Z_{i}^{\mathrm{Q}}\}_{i=1}^{n_{\mathrm{Q}}}$ where $Z_{i}^{\mathrm{Q}}=(X_{i}^{\mathrm{Q}},\pi^{\mathrm{Q}}(X_{i}^{\mathrm{Q}}),Y_{i}^{\mathrm{Q},(\pi^{\mathrm{Q}}(X_{i}^{\mathrm{Q}}))}).$
  • For each of the four methods (i.e. LDPMAB, Linear, NN, ABSE), we now start the learning process on the target data for $i\in[n_{\mathrm{P}}]$, where note that given the pulled arm $\pi_{i}(X_{i}^{\mathrm{P}})$, the reward is generated via $Y_{i}^{\mathrm{P},(\pi_{i}(X_{i}^{\mathrm{P}}))}:=\mathbf{1}(\dot{Y}_{i}^{\mathrm{P}}=\pi_{i}(X_{i}^{\mathrm{P}}))$. The cumulative reward is therefore $\sum_{i=1}^{n_{\mathrm{P}}}{Y}_{i}^{\pi_{i}(X_{i})}$.
Note that all three steps above involves randomness, stemming from permutation, realization of behavior policy, the privacy mechanism (i.e. Laplacian random noises), and realization of target policy.
The experiment results for each method (LDPMAB, Linear, NN) on the three datasets are summarized in Table 2 under various combinations of privacy budgets $(\varepsilon,\varepsilon_{m})$. Note that to standardize the scale across datasets, we report the ratio of the mean reward of each method relative to that of ABSE, which, as discussed above, is implemented on the target data non-privately without transfer learning. Thus, a reported value larger than $1$ means that the method is better than ABSE and vice versa.
@@
Several observations are in order. First, LDPMAB with auxiliary data outperforms its competitors in terms of both best performance (number of significantly better rewards) and average performance (rank-sum). This shows that our proposed methods can effectively utilize auxiliary data and thus achieves knowledge transfer with the designed jump-start scheme. In contrast, Linear and NN occasionally have negative transfer, where auxiliary data worsens the performance. In addition, without auxiliary data, LDPMAB still outperforms Linear, suggesting the advantage of the nonparametric nature of LDPMAB. Compared to ABSE, the competing methods without auxiliary data are usually worse (i.e. with ratio less than 1) since LDP is required, indicating the cost of privacy.
@@
Table 2: The best performer among 6 methods (i.e. LDPMAB w/wo, Linear w/wo, NN w/wo) are marked in bold for each dataset under different combinations of $(\varepsilon,\varepsilon_{m})$. Note that for each dataset, we report the performance at both $t=n_{\mathrm{P}}/4$ and $t=n_{\mathrm{P}}$ to highlight the effect of transfer learning. To ensure statistical significance, we adopt the Wilcoxon signed-rank test (Wilcoxon,, 1992) with a significance level of 0.05 to check if the result is significantly better. The best results that hold significance towards the others are highlighted in grey.
@@ 5 Conclusions and Discussions
In this work, we investigate the problem of nonparametric contextual multi-armed bandits under local differential privacy. We propose a novel uniform-confidence-bound based algorithm, which achieves near-optimal performance supported by a newly derived minimax lower bound. To further improve the performance limit of LDP contextual MAB, we consider transfer learning, which incorporate side information from auxiliary datasets that are also subject to LDP constraints. Assuming covariate shift, we introduce a jump-start scheme to leverage the auxiliary data, attaining the established minimax lower bound, up to logarithmic factors in interesting regimes. Extensive experiments on synthetic and real datasets validate our theoretical findings and demonstrate the superiority of our methodology.
We remark on the implications of our method in the context of multi-task learning. Consider a scenario where a set of $M$ players are deployed to engage in a bandit game, with the overall objective being to minimize the average regret across all players (Deshmukh et al.,, 2017; Wang et al.,, 2021). These players simultaneously interact with a shared set of arms. At each round, each player selects an arm and receives feedback. The conditional distribution of each arm’s reward is identical across all players. Under this setting, the estimator in (26) is permutation invariant with respect to the datasets. This means that treating any dataset as the target dataset does not affect the estimator’s effectiveness or the subsequent confidence bound (28). This observation suggests that the proposed methodology can be extended to multi-task learning, provided Algorithm 3 is adapted to accommodate parallel interactions. We leave a thorough investigation for future research.